OpenStudy (anonymous):

HELP PLEASE A skydiver of total mass 100kg leaves an aircraft flying at 3000m and freefalls down to a height of 1200m. By this time, she has reached a constant vertical velocity of 90m/s. (i)Calculate the loss of gravitational potential energy during the fall. (ii)Calculate the gain in kinetic energy of the skydiver (iii) Hence calculate the mean air resistance force acting on the skydiver during the fall.

OpenStudy (anonymous):

(i) The skydiver fall 1800 meters before reaching terminal velocity. The change in potential energy is: m g delta h = 100 kg * 9.81 m/s^2 * 1800 m = 1,765,800 Joules (ii) ignoring any horizontal speed (perhaps the skydiver dropped from a balloon :) the starting speed was zero. The final speed was 90 m/s so Ke = (1/2) m v^2 = 50*8100 = 405,000 Joules (iii) So the work done by air friction was 1,765,800 - 405,000 = 1,360,800 Joules Force * distance = work done F * 1800 meters = 1,360,800 Joules F = 756 Newtons